Home Up

1998 Summary

1998 WODC Trail Work Summary

Please see the photos of the 1998 ice storm.


Trail


Work Done


Current Condition

Work Hours (including hiking time but not travel)

Trips

Crew

Adopter

Total

Bennett St Cleared major storm damage and some drainages. (Also cleared much of Bennett S. Trl on one trip.) Many drainages require clearing. Some erosion damage & recent blowdowns.

270

8.5

incl

?

279

Bickford Major storm damage cleared, mostly by adopter David Bowles. Drainage work (water bars) needed on descent to Paugus Mill.

0

10

12

?

22

Big Rock Cave Storm damage and drainages cleared mostly by Over The Hill Hikers. Trail in generally good condition.

0

8

165

173

B. Ledge Light to moderate storm damage cleared. Drainages cleared. Generally good condition, but needing further erosion control work and stepping stones.

75

36

?

111

B.L. Cutoff Moderate storm damage cleared. Drainages cleared. Needs erosion control work.

49

9.25

2

?

60

Brook Path Light storm damage cleared. Good condition except for ongoing change of river channel and seasonal bridge washout.

0

0

2

?

2

Cabin Moderate storm damage cleared. Considerable drainage work by Light Crew. Brushed. Good condition. Muddy sections need ongoing work.

41

62.75

6

?

110

Dicey's Mill Heavy storm damage fully cleared. All 200+ drainages cleared. Brushed. Very good condition. Some erosion problems below Wiggin.

150.5

19.5

?

incl

170

East Loop Light storm damage cleared. Good condition except for steep descent to Walden Trl

0

4

?

4

Gordon Path Blazing and signs restored. No storm damage. Good condition. Needs marker posts across field at East end.

13

0

?

13

Kate Sleeper Moderate storm damage cleared. Well brushed. Drainages cleared. Considerable drainage improvements. Good condition. Drainage & erosion work required in 3/4 mile below Camp Shehadi.

392

0

-

392

Kelley Heavy storm damage cleared, including several uprooted trees. Brushed. Major treadway construction and repair completed. Good condition. Recent erosion damage in top 1/2 mile must be monitored.

249

54

incl

302

Lawrence Heavy storm damage cleared. Brushed. Poor to good condition. Major erosion problems at Overhang and at wooden steps on Paugus.

61

25.5

16

103

McCrillis Path Light storm damage cleared. Brushed. Drainages cleared. Very good within WMNF. Significant erosion, blazing, and signage problems outside WMNF.

0

15.75

?

16

Old Mast Rd Heavy storm damage cleared. Drainages cleared. Many new water bars installed to halt rain damage. Campsites on Smalley/Boettiger land closed and naturalized in cooperation with landowner. Gravel pit and dump to be closed by landowner. Very good condition. Recent erosion must be monitored. Signs of increased Mtn bike use. "No Camping" areas will require ongoing attention.

113

110.25

?

223

Pasture P. Light to moderate storm damage cleared. "Trail->" sign placed at Mt. Katherine terminus. Good condition. Rich Forest sign needs to be rehung. Needs reblazing.

0

8

?

8

Red Path Light storm damage cleared. New signs placed in cooperation with Chandlers. Reblazed. Good condition. New water bars needed on ascent to Tilton Spring.

27

0

incl

?

27

Rollins Light storm damage cleared. Good condition. Drainage work needed on East end.

0

0

20

20

Short Cut Moderate storm damage cleared. Good condition. Needs reblazing.

0

4

?

4

Sq. Ledge Moderate storm damage cleared. Brushed. Drainages cleared. New drainage and erosion-control work accomplished. New stepping stones. Two sections moved from drainages. Generally good condition. Short relos must be monitored. Some confusion regarding Peregrine signage.

186

125

?

incl

311

Tilton Spr. Light storm damage cleared. Good condition. Needs reblazing.

9

8

?

17

Tom Wiggin Moderate storm damage cleared. Drainage work before stream. Poor condition. Significant erosion on steep sections will require major trail work. Relocation or closure must also be considered.

28

39.25

?

67

Walden Light storm damage cleared. Funding has been obtained for complete treadway restoration. Poor to fair condition. No success to date in securing experienced trail crew.

44

29.5

?

74

Won. Range Light to moderate storm damage cleared. Good condition. Treadway work needed by brook.

30

14.5

?

45

*** USFS & CMC trails The Light Crew also cleared these USFS & CMC trails: McCrillis Trail, Old Paugus Trail, Bee Line (west), Bee Line Cutoff, Olivarian Brook (in Wilderness), Sq. Ledge Branch, Whittin Brook Trl.

10

155.5

-

166

*** Camp Heermance No work done this year other than routine monitoring. Poor condition. Shelter is structurally unsafe, with part of north wall missing. Privy is full, seat is rusted, and deck is rotting.

-

-

-

0

*** Camp Shehadi No work done this year other than routine monitoring. Poor condition. Shelter continues to settle into ground. Headroom is minimal. Earth floor is often wet.

-

-

-

0

*** Camp Rich Several campsites near spring and trail were removed. Pick up trash. Close heard paths. Clean fire ring. Shelter in fair condition. (Roof leaks.) Outhouse in good condition. Viable tenting options nearby. Excellent water supply.

18.5

18.5

0

37

 

Totals Work Hours

1,765

766

223

2,754

 

Send comments about this web site to      This page was last modified on 03/07/2007 .